OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2). −9 −7 −10 7 @Nurali

OpenStudy (nurali):

f(x) = −4x − 1 f(−2)=-4(-2)-1 f(−2)=8-1 f(−2)=______

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

@Nurali